OpenStudy (anonymous):

I need help with this problem..do I mutiply or divide the number in parthensis? 19(24) = 19(20)+19(4)

OpenStudy (anonymous):

in math, a number in parenthesis is multiplied. So this would be 19*24 = 19*20 + 19*4 Which is true and the rational behind the distributive property of multiplication.

OpenStudy (anonymous):

okay..so 456=380 + 76 = 456? is this correct?

OpenStudy (anonymous):

The equation is true, but that could have been trivially confirmed with a calculator.

OpenStudy (anonymous):

but am i on the right track?

OpenStudy (anonymous):

I don't understand what you mean. a*(b+c) = a*b + a*c by the distributive rule of multiplication

OpenStudy (anonymous):

i reaaly don't get what u just wrote...but thanks.. just a 6th grader needing help with a problem

OpenStudy (anonymous):

I wrote the general form for the principle you were asking about. Just substitute a, b, and c for whatever numbers you like.

OpenStudy (anonymous):

e.g. a=2, b=3, c=4 2*(3+4) = 2*7 = 14 2*3 + 2*4 = 6 + 8 = 14